OpenStudy (anonymous):

Solve for x. 11x + 4 < 37 x < 37/15 x < 3 x < 41/11 x < 363

OpenStudy (shelld93):

the aim is to get x on its own. remember that what you do to one side you must do to the other

OpenStudy (shelld93):

do you have any idea ?

OpenStudy (anonymous):

no not really

OpenStudy (shelld93):

okay so your first step would be th remove the +4 on the right hand side. to do this you minus 4 from both sides. the equation should now look like this: 11x + 4 - 4 = 37 - 4

OpenStudy (anonymous):

ok

OpenStudy (shelld93):

of course 4 - 4 = 0 so not the +4 on the right hand side is cancelled out. 11x = 37-4

OpenStudy (anonymous):

answer is 3

OpenStudy (shelld93):

yeah it is.

OpenStudy (anonymous):

dont forget to close this question!

OpenStudy (anonymous):

thanks guys my bad I had to edit the question